# Large Language Models as Natural Selector for Embodied Soft Robot Design

### v1

@@ Abstract
Designing soft robots is a complex and iterative process that demands cross-disciplinary expertise in materials science, mechanics, and control, often relying on intuition and extensive experimentation. While Large Language Models (LLMs) have demonstrated impressive reasoning abilities, their capacity to learn and apply embodied design principles—crucial for creating functional robotic systems—remains largely unexplored. This paper introduces RoboCrafter-QA, a novel benchmark to evaluate whether LLMs can learn representations of soft robot designs that effectively bridge the gap between high-level task descriptions and low-level morphological and material choices. RoboCrafter-QA leverages the EvoGym simulator to generate a diverse set of soft robot design challenges, spanning robotic locomotion, manipulation, and balancing tasks. Our experiments with state-of-the-art multi-modal LLMs reveal that while these models exhibit promising capabilities in learning design representations, they struggle with fine-grained distinctions between designs with subtle performance differences. We further demonstrate the practical utility of LLMs for robot design initialization. Our code and benchmark will be available to encourage the community to foster this exciting research direction1.
@@ I Introduction
Figure 1 illustrates a conceptual shift in the paradigms of creature and robot design. While biological evolution (Fig. 1a) and human-engineered design (Fig. 1b) have driven progress, they are inherently slow and limited by human cognitive capacity. The emergence of Artificial Intelligence, particularly Large Language Models (LLMs), presents a new opportunity: LLMs as a form of ‘natural selector’ for robot design (Fig. 1c). LLMs, trained on vast datasets that encompass text and code, have demonstrated impressive capabilities in the understanding, generation, and even reasoning of natural languages. As discuss in [5], LLMs have the potential to revolutionize the robotic design process by serving as intelligent co-designers, however, there is no work that enables comprehensive and quantitative analyses and understanding of LLM for robot design. This raises a fundamental question: Can LLMs leverage their latent representations for the design of embodied soft robots, bridging the gap between high-level task descriptions and low-level morphological choices?
@@
Figure 1: Conceptual comparison of creature/robot evolution paradigms. (a) Evolution-driven emergence of creatures in nature through selective pressures; (b) Traditional human-engineered robot design guided by intuition and expertise; (c) AI-driven robot design, where large language models (LLMs) act as a ‘natural selectorin the evolution of soft robotics, e.g., modular robot shown in the figure. This shift highlights the transition from biological evolution to human-driven engineering and finally to AI-mediated selection.
Existing LLM benchmarks focus primarily on general-purpose language understanding, code generation, or visual reasoning in non-embodied contexts. Although these benchmarks are valuable, they do not adequately assess the ability of LLMs to reason about physical systems, material properties, and the complex interplay between a robot’s morphology, its environment, and the task it must perform. In essence, current LLM benchmarks do not test the embodied design intelligence required for successful soft robot creation. The crucial question of whether LLMs can effectively learn and apply design principles for embodied soft robots remains largely unanswered.
To address this critical gap, we introduce RoboCrafter-QA, a novel benchmark specifically designed to evaluate the understanding of embodied soft robot design of multimodal LLMs. The entire pipeline, encompassing data generation, question formulation, and LLMs evaluation, is illustrated in Figure 2. RoboCrafter-QA leverages the EvoGym robotic simulator [6], a high-fidelity, computationally efficient platform for evolutionary robotics research. EvoGym allows us to generate a large-scale, controlled dataset of (environment, task, soft robot design, performance) tuples, providing a rich ground truth for evaluating soft robot design understanding. The core of RoboCrafter-QA is a challenging Question-Answering (QA) task: given a textual description of a simulated environment and task, a LLM must select the better-performing soft robot design from a pair of candidates. This task directly probes the LLM’s ability to learn and apply representations of embodied design principles, requiring it to reason about the intricate interplay of environmental factors, task objectives, and the morphological and material properties of the robot designs. Success hinges not on simple pattern matching, but on a deeper understanding of how design choices impact performance in a physically realistic simulation.
@@
Figure 2: RobotCrafter-QA data generation and LLM evaluation pipeline.
@@
Our benchmarking results, using state-of-the-art LLMs, reveal both the potential and the limitations of current models in this challenging domain. We find that while LLMs can learn to distinguish between significantly different designs, they struggle with fine-grained distinctions, highlighting the need for further advancements in representation learning for embodied design. Finally, beyond leveraging LLMs as robot design selectors, we show that they can also aid in refining and optimizing initial designs by incorporating task-specific constraints and physical reasoning.
@@
    We introduce RoboCrafter-QA, a novel benchmark for evaluating the embodied soft robot design understanding of multi-modal LLMs, focusing on the evaluation of design representations.
@@
    We propose a challenging Question-Answering task that directly assesses LLMs’ ability to reason about the interplay between task requirements, environment constraints, and soft robot design, going beyond traditional pattern recognition.
@@
    We provide a comprehensive evaluation of state-of-the-art multi-modal LLMs on the RoboCrafter-QA benchmark, offering insights into their capabilities and limitations in applying embodied design principles.
@@
    We showcase that LLMs generate more effective robot design initializations than random initializations when provided with a textual task description.
@@ II-A Soft Modular Robot Design and Optimization
Soft modular robots exploit distributed actuation and compliance to enhance flexibility and adaptability in dynamic environments [7, 8, 9, 10]. Unlike rigid robots with fixed kinematics, these systems dynamically alter morphology and stiffness using soft actuation mechanisms such as pneumatic actuators, shape memory alloy, and electroactive polymers. Recent studies have emphasized reconfigurability as a key advantage. Zhang et al. [11] introduced a lattice-based framework for morphology adaptation under external forces, while Li and Banerjee [12] developed modular designs for flexible attachment and detachment. Further approaches integrate sensor feedback with adaptive materials to enable real-time reconfiguration [13, 14]. However, optimizing soft modular robots remains challenging due to nonlinear deformations and complex material interactions. Traditional techniques, such as evolutionary algorithms and reinforcement learning [15, 16], provide heuristic-based solutions but struggle with scalability. Recent advances leverage generative AI for automated morphology discovery, with DiffuseBot [17] integrating diffusion-based generative modeling and differentiable physics, while Garcia et al. [18] combine data-driven and physics-based constraints for enhanced design efficiency.
Our work advances this field by benchmarking LLMs for automated robot design and proposing methods to systematically probe their reasoning in morphology optimization. Leveraging the EvoGym simulator as a high-fidelity testbed, we enable structured evaluations of LLM-driven design selection and optimization under realistic physics constraints.
@@ II-B LLM Evaluation Benchmarks
General-purpose benchmarks such as BIG-Bench [19], MMLU [20], and C-Eval [21] assess LLM performance on linguistic fluency, knowledge retrieval, and reasoning. However, the rapid evolution of LLMs has led to specialized evaluations like ScienceQA [22, 23, 24] and MMLU-pro [25, 26], which test domain-specific reasoning in scientific and professional contexts. These benchmarks provide deeper insights into specialized knowledge but remain primarily focused on textual understanding. While these benchmarks evaluate language comprehension, they often overlook embodied intelligence—where physical constraints, environmental interactions, and morphology influence task success. Recent research has emphasized the need for evaluations that integrate physical reasoning, as LLMs often struggle with robotics-related decision-making [27]. Studies highlight the importance of specialized assessments that bridge high-level planning with real-world constraints [28, 29]. Without such evaluations, LLMs may demonstrate strong linguistic capabilities but fail to generalize effectively in real-world applications requiring physical interactions.
Existing benchmarks focus on textual reasoning but fail to measure how well LLMs incorporate task goals, environmental conditions, and morphological factors in robotics. To address this gap, our proposed RoboCrafter-QA benchmark provides a structured framework to assess how effectively LLMs reason about soft robot design choices. By incorporating multimodal constraints and physical reasoning, RoboCrafter-QA extends beyond conventional textual benchmarks, offering a more comprehensive evaluation of embodied intelligence in AI systems. Furthermore, it facilitates a deeper understanding of how LLMs navigate the intersection of design trade-offs, material properties, and task-specific constraints in robotics applications.
@@ III-A Benchmark Overview
The RoboCrafter-QA benchmark is designed to rigorously evaluate the embodied design reasoning capabilities of LLMs in the context of soft robot design over 12 tasks as illustrated in Figure 3. It comprises a curated dataset of Question-Answer (QA) pairs, $\{(Q_{i},G_{i})\}_{i=1}^{N}$, generated using the EvoGym simulator. Each question $Q_{i}=(D_{Q_{i}},R_{1_{i}},R_{2_{i}})$ presents a LLM with a textual prompt $D_{Q_{i}}$ and two candidate robot designs, $R_{1_{i}}$ and $R_{2_{i}}$, represented as voxel-based material matrices. The ground truth answer, $G_{i}\in\{A,B\}$, indicates the robot design that achieves the higher reward, $P(R,\pi,E)$, in the simulated environment $E$. The benchmark encompasses a diverse range of task environments, $E=(D,T,C)$, including locomotion, object manipulation, and climbing, each characterized by a textual description $D$, a task objective $T$, and environmental constraints $C$. An example of a question $Q_{i}$ and its components is illustrated in Figure 4. The difficulty of each question is modulated by the difference in performance rewards, $|P(R_{1_{i}},\pi,E)-P(R_{2_{i}},\pi,E)|$, between the candidate robot designs, enabling a nuanced assessment of LLMs’ ability to discern subtle design trade-offs. An example of various robot designs and their corresponding rewards for a pushing task and a bridge walking task is shown in Figure 5. The dataset is partitioned into Easy and Hard subsets based on this performance difference, providing a comprehensive evaluation of LLMs’ embodied design reasoning across varied complexity.
Figure 3: Robot design tasks in our RoboCrafter-QA benchmark.
@@ III-C Question Generation
Context Integration. Task descriptions $T$ and difficulty levels are retrieved from predefined dictionaries. Each question includes environment-specific objectives, such as ”Climb as high as possible on a vertical wall” or ”Walk as far as possible on flat terrain” to provide necessary context $D_{Q_{i}}$.
@@ III-D Benchmark Statistics
Table I provides a detailed overview of the RoboCrafter-QA benchmark dataset, showcasing its breadth and controlled difficulty across twelve distinct EvoGym tasks. Comprising a substantial 36,000 questions in total, with 3,000 questions dedicated to each environment, the benchmark ensures a robust evaluation. The ’Reward Diff’ column highlights a critical aspect of question difficulty, presenting the range of performance score differences between the paired robot designs. This range varies across environments, from subtle distinctions (e.g., Climber-v0 with 0.01 to 0.19) to more pronounced performance gaps (e.g., Jumper-v0 with 0.13 to 8.07), effectively modulating the challenge for LLMs. Furthermore, the ’Max Reward’ and ’Min Reward’ columns contextualize the performance landscape within these environments, demonstrating the variability in achievable and poor robot designs that contribute to the question generation process. This statistical breakdown underscores the dataset’s comprehensive nature, designed to rigorously challenge and differentiate the embodied design reasoning capabilities of multi-modal LLMs across a spectrum of task complexities and performance disparities.
@@ IV Experiments
IV-A Evaluation Metrics
Accuracy: Measures the proportion of model-generated responses that match the expected answers. Given $N$ test queries, accuracy $A$ is computed as:
$A=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(O_{i}=G_{i})$ (1)
where $O_{i}$ is the model’s output, $G_{i}$ is the ground truth, and $\mathbb{I}(\cdot)$ is an indicator function returning 1 if responses match, 0 otherwise.
Consistency: Evaluates the stability of a large language model by measuring the reproducibility of the response in multiple trials. Each query is presented three times, and the consistency score $\kappa$ is defined as:
$\kappa=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbb{I}(O_{i}^{(1)}=O_{i}^{(2)})+\mathbb{I}(O_{i}^{(1)}=O_{i}^{(3)})}{2}$ (2)

### v2

# Natural Selection via Foundation Models for Soft Robot Evolution

@@ Abstract
Designing soft robots is a complex and iterative process that demands cross-disciplinary expertise in materials science, mechanics, and control, often relying on intuition and extensive experimentation. While foundation models, especially Large Language Models (LLMs), have demonstrated impressive reasoning abilities, their capacity to conduct embodied design remains largely unexplored. This paper introduces RoboCrafter-QA, a novel benchmark to evaluate whether LLMs can learn representations of soft robot designs that effectively bridge the gap between high-level task descriptions and low-level morphological and material choices. RoboCrafter-QA leverages the EvoGym simulator to generate a diverse set of soft robot design challenges, spanning robotic locomotion, manipulation, and balancing tasks. Our experiments with SOTA multi-modal LLMs reveal that while these models exhibit promising capabilities in learning design representations, they struggle with fine-grained distinctions between designs with subtle performance differences. To overcome these limitations, we finetune an efficient, open-source LLM that achieves SOTA performance on our benchmark, demonstrating superior capabilities in both design selection and direct generation of high-performing robot morphologies. Furthermore, we construct a physical replica of the modular soft robot and demonstrate a strong sim-to-real correlation, validating that superior benchmark performance has the potential to translate to effective real-world design selection. Our full system will be open-sourced to foster this exciting direction.
@@ I Introduction
Fig. 1 illustrates a conceptual shift in the paradigms of creature and robot design. While biological evolution (Fig. 1a) and human-engineered design (Fig. 1b) have driven progress, they are inherently slow and limited by human cognitive capacity. The emergence of foundation models, particularly Large Language Models (LLMs), presents a new opportunity: LLMs as a form of ‘natural selector’ for robot design (Fig. 1c). LLMs, trained on vast datasets that encompass text and code, have demonstrated impressive capabilities in the understanding, generation, and even reasoning of natural languages. As discussed in [5], LLMs have the potential to revolutionize the robotic design process by serving as intelligent co-designers, however, a comprehensive and quantitative analyses and understanding of LLM for robot design is still underexplored. This raises our fundamental research question: Can LLMs leverage their latent understanding for the design of embodied soft robots, bridging the gap between high-level task descriptions and low-level morphological choices?
@@
Figure 1: Conceptual comparison of natural creature and artificial robot evolution paradigms. (a) Evolution-driven emergence of creatures in nature through selective pressures; (b) Traditional human-engineered robot design guided by intuition and expertise; (c) AI-driven robot design selection, where foundation models (e.g., Large Language Models) act as a ‘natural selector’ for robot design evolution (e.g., modular robots shown in the figure). This shift highlights the transition from biological evolution to human-driven engineering and finally to AI-empowered selection.
@@
Figure 2: RobotCrafter-QA robot design data generation and LLM evaluation pipeline.
Existing LLM benchmarks are ill-suited for this problem. They focus on general-purpose language, code, or non-embodied visual reasoning [6, 7, 8, 9], failing to assess an LLM’s ability to reason about physical systems. Critically, they do not test the embodied design intelligence required for soft robot creation—namely, the complex interplay between a robot’s morphology, its environment, and the task it must perform. To address this gap, we introduce RoboCrafter-QA, a benchmark for evaluating the embodied design intelligence of multimodal LLMs (pipeline in Fig. 2). The benchmark is built using the EvoGym simulator [10] to generate a large-scale dataset of (environment, task, soft robot design, performance) tuples. The core task is a Question-Answer (QA) challenge: given a textual description of a task and environment, an LLM must select the better-performing robot design from a pair. This format directly tests the LLM’s ability to reason about the complex interplay between morphology, materials, and performance, moving beyond simple pattern matching.
@@
Our benchmarking results, using state-of-the-art LLMs, reveal both the potential and the limitations of current models in this challenging domain. We find that while LLMs can learn to distinguish between significantly different designs, they struggle with fine-grained distinctions, highlighting the need for further advancements in representation learning for embodied design. To push beyond these limitations, we demonstrate that an efficient, open-source LLM can be finetuned on RoboCrafter-QA to achieve state-of-the-art design selection. We show this model not only surpasses SOTA models in the selection task but can also be prompted to generate novel, high-performing robot morphologies directly. Finally, to close the loop and validate the practical utility of our benchmark, we construct a physical modular robot system. Our real-world experiments confirm a strong sim-to-real correlation, demonstrating that the design principles learned by our model in simulation translate to effective performance in a physical robot.
@@
    We introduce RoboCrafter-QA, a novel multi-modal question-answering benchmark that assesses an LLM’s ability to reason about the interplay between task requirements and soft robot design.
@@
    We provide a comprehensive evaluation of state-of-the-art M-LLMs on RoboCrafter-QA, revealing key insights into their current capabilities and limitations in embodied design reasoning.
@@
    We finetune an efficient, open-source M-LLM that establishes a new state-of-the-art on our benchmark, demonstrating significantly stronger robot design selection and reasoning capabilities.
@@
    We construct a physical replica of the modular soft robot and demonstrate a strong sim-to-real correlation, validating that superior benchmark performance translates to effective real-world design selection.
Figure 3: Overview of robot design tasks curated from EvoGym [10] for the RoboCrafter-QA benchmark.
@@ II-A Soft Modular Robot Design and Optimization
Soft modular robots leverage compliance and distributed actuation (e.g., pneumatics [11], SMAs [12], EAPs [13]) to achieve adaptability and dynamic morphology, unlike their rigid counterparts [14, 15, 16, 17]. Research has focused on reconfigurability, including lattice-based adaptation to external forces [18], flexible attachment mechanisms [19], and real-time, sensor-driven reconfiguration [20, 21].
Optimizing these robots is challenging due to their nonlinear dynamics. Early approaches pioneered evolutionary algorithms for automatic design [22], a method later used to create xenobots from biological cells [23]. More recent work leverages generative AI, such as diffusion models with differentiable physics [24] and hybrid data-driven/physics-based methods [25], for automated morphology discovery. Our work contributes by benchmarking LLMs for automated design selection, using the EvoGym simulator to systematically probe their reasoning on morphology optimization under realistic physical constraints.
@@ II-B LLM Evaluation Benchmarks
General benchmarks like MMLU [26] and C-Eval [27] test broad linguistic and reasoning skills. In response to rapid model evolution, specialized benchmarks such as ScienceQA [7, 8] and MMLU-pro [28, 29] have emerged to test deep, domain-specific knowledge, though they remain focused on textual understanding.
These text-centric benchmarks overlook embodied intelligence, where physical constraints and morphology are critical. LLMs are known to struggle with robotics-related physical reasoning [30], highlighting the need for specialized assessments that bridge high-level planning with real-world physical constraints [31, 32]. Existing benchmarks fail to measure how LLMs integrate task goals, environmental conditions, and morphology. We introduce RoboCrafter-QA, a structured benchmark to assess how LLMs reason about soft robot design choices. RoboCrafter-QA moves beyond purely textual evaluation to provide a comprehensive assessment of embodied intelligence.
@@ III-A Benchmark Overview
The RoboCrafter-QA benchmark is designed to rigorously evaluate the embodied design reasoning capabilities of LLMs in the context of soft robot design over 12 tasks as illustrated in Fig. 3. It comprises a curated dataset of Question-Answer (QA) pairs, $\{(Q_{i},G_{i})\}_{i=1}^{N}$, generated using the EvoGym simulator. Each question $Q_{i}=(D_{Q_{i}},R_{1_{i}},R_{2_{i}})$ presents a LLM with a textual prompt $D_{Q_{i}}$ and two candidate robot designs, $R_{1_{i}}$ and $R_{2_{i}}$, represented as voxel-based material matrices. The ground truth answer, $G_{i}\in\{A,B\}$, indicates the robot design that achieves the higher reward, $P(R,\pi,E)$, in the simulated environment $E$. The benchmark encompasses a diverse range of task environments, $E=(D,T,C)$, including locomotion, object manipulation, and climbing, each characterized by a textual description $D$, a task objective $T$, and environmental constraints $C$. An example of a question $Q_{i}$ and its components is illustrated in Fig. 4. The difficulty of each question is modulated by the difference in performance rewards, $|P(R_{1_{i}},\pi,E)-P(R_{2_{i}},\pi,E)|$, between the candidate robot designs, enabling a nuanced assessment of LLMs’ ability to discern subtle design trade-offs. An example of various robot designs and their corresponding rewards for a pushing task and a bridge walking task is shown in Fig. 5. The dataset is partitioned into Easy and Hard subsets based on this performance difference, providing a comprehensive evaluation of LLMs’ embodied design reasoning across varied complexity.
@@ III-B Data Generation
Task Name # Questions Reward Diff Max Reward Min Reward
Walker-v0 3,000 0.74 to 2.89 4.09 -1.21
BridgeWalker-v0 0.39 to 3.69 4.04 -1.49
BidirectionalWalker-v0 0.31 to 4.16 1.81 -3.69
UpStepper-v0 0.08 to 2.51 1.64 -1.27
GapJumper-v0 0.08 to 2.62 2.28 -0.77
Carrier-v0 0.53 to 3.50 2.81 -2.86
Carrier-v1 0.90 to 6.62 3.60 -5.85
Pusher-v0 0.49 to 3.07 3.07 -1.66
Pusher-v1 0.04 to 1.39 0.21 -1.33
Climber-v0 0.01 to 0.19 0.16 -0.18
Jumper-v0 0.13 to 8.07 0.04 -8.44
Balancer-v0 0.12 to 1.05 0.13 -1.41
Summary 36,000 0.01 to 8.07 4.09 -8.44
@@ III-C Question Generation
Context Integration. Each question includes environment-specific objectives, such as ‘Climb as high as possible on a vertical wall’ or ‘Walk as far as possible on flat terrain’ to provide necessary context.
@@
Category Setting Heuristic Qwen-0.5-Instruct GPT5 Gemini-2.5-Pro Grok-4
Accuracy Accuracy Consistency Accuracy Consistency Accuracy Consistency Accuracy Consistency
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Locomotion Walker-v0 49.61 54.03 39.17 37.83 39.00 38.33 51.34 49.34 76.67 78.67 48.33 49.84 66.67 69.67 60.33 56.84 72.00 70.33
BridgeWalker-v0 45.53 46.59 37.67 38.67 40.00 42.67 54.33 52.84 70.67 71.67 61.16 52.33 71.67 66.00 69.33 56.66 74.67 68.00
BidirectionalWalker-v0 48.79 51.91 36.33 36.50 37.33 38.00 50.66 53.84 73.33 77.67 51.66 47.50 70.00 69.67 48.50 51.50 76.33 77.00
UpStepper-v0 57.14 53.06 41.83 41.17 48.33 41.00 67.17 51.66 82.33 85.33 61.66 51.66 69.33 68.67 64.83 56.50 78.33 76.33
GapJumper-v0 51.29 50.66 37.67 38.17 35.67 36.33 58.00 51.84 72.67 74.33 60.17 57.17 65.67 67.00 64.67 57.66 68.67 66.67
Average 50.47 51.25 38.53 38.47 40.07 39.27 56.30 51.90 75.13 77.53 56.60 51.70 68.67 68.20 61.53 55.83 74.00 71.67
Object Manipulation Carrier-v0 45.07 51.02 38.67 37.17 33.33 31.33 69.16 57.84 82.33 79.67 60.34 51.83 68.00 70.33 68.34 58.34 77.33 72.00
Carrier-v1 40.22 43.66 34.50 35.00 40.67 36.00 65.34 59.83 81.33 75.67 67.16 61.33 77.00 74.67 65.00 57.84 72.67 71.00
Pusher-v0 45.26 57.19 38.33 41.00 42.67 42.33 30.34 45.84 74.00 76.33 40.50 50.66 73.67 76.00 58.50 59.66 69.00 72.67
Pusher-v1 56.01 52.30 39.33 38.50 40.67 39.33 49.16 48.84 72.33 79.00 51.00 54.67 68.00 74.00 53.84 56.50 76.33 81.00
Average 46.64 51.04 37.71 37.92 39.33 37.25 53.50 53.09 77.50 77.67 54.75 54.62 71.67 73.75 61.42 58.09 73.83 74.17
Climbing & Balancing Climber-v0 54.04 55.82 37.00 37.00 37.67 36.33 51.50 49.84 79.00 75.67 55.50 57.00 71.67 75.33 57.34 57.67 79.33 77.33
Jumper-v0 18.13 33.42 36.00 37.00 42.67 41.00 9.84 31.50 91.67 82.33 16.84 39.67 86.33 78.67 42.16 44.34 72.33 68.00
Balancer-v0 53.71 58.38 34.83 36.50 33.00 37.67 54.66 52.33 70.00 72.67 51.17 59.16 68.33 69.67 63.84 57.34 76.33 68.00
Average 41.96 49.21 35.94 36.83 37.78 38.33 38.67 44.56 80.22 76.89 41.17 51.94 75.44 74.56 54.45 53.12 76.00 71.11
Average 47.07 50.67 37.61 37.88 39.25 38.36 50.96 50.46 77.19 77.42 52.12 52.73 71.36 71.64 59.72 55.90 74.44 72.36
@@ III-D Benchmark Statistics
Table I provides a detailed overview of the RoboCrafter-QA benchmark dataset, showcasing its breadth and controlled difficulty across twelve distinct EvoGym tasks. Comprising a substantial 36,000 questions in total, with 3,000 questions dedicated to each environment, the benchmark ensures a robust evaluation. The ‘Reward Diff’ column highlights a critical aspect of question difficulty, presenting the range of performance score differences between the paired robot designs. This range varies across environments, from subtle distinctions (e.g., Climber-v0 with 0.01 to 0.19) to more pronounced performance gaps (e.g., Jumper-v0 with 0.13 to 8.07), effectively modulating the challenge for LLMs. Furthermore, the ‘Max Reward’ and ‘Min Reward’ columns contextualize the performance landscape within these environments, demonstrating the variability in achievable and poor robot designs that contribute to the question generation process. This statistical breakdown underscores the dataset’s comprehensive nature, designed to rigorously challenge and differentiate the embodied design reasoning capabilities of multi-modal LLMs across a spectrum of task complexities and performance disparities.
@@ IV Experiments
IV-A Evaluation Metrics of Robot Design Selection
Accuracy. It measures the proportion of model-generated responses that match the expected answers. Given $N$ test queries, accuracy $Acc$ is computed as:
$Acc=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(O_{i}=G_{i})$ (1)
where $O_{i}$ is the prediction, $G_{i}$ is the ground truth, and $\mathbb{I}(\cdot)$ is an indicator function returning 1 if responses match, 0 otherwise.
Consistency. It evaluates the stability of a model by measuring the reproducibility of the response in multiple trials. Each query is presented three times, and the consistency is:
$Cons=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbb{I}(O_{i}^{(1)}=O_{i}^{(2)})+\mathbb{I}(O_{i}^{(1)}=O_{i}^{(3)})}{2}$ (2)